(Laplace Transforms) Find a function whose Laplace transform is 1 $3(s – 2)
Answer #1

- Solution - het FIS) = 5215-2) FIS) = is convert isto partial fractions 5215-2) А B C 0 + 52(S-2) S2 S-2 = AS (S-2) +B15-2)comparing a2 coefficients, A +(= 0 At C(=174) 4 ile Substituting these values in equation © 4 Fis)e 3.0 s-2 1 FIS) = -1 45 45